Income Statement Pietro Frozen Foods, Inc., produces frozen pizzas. For next year, Pietro predicts that 49,000 units will be produced, with the following total costs: Direct materials Direct labor Variable overhead Fixed overhead Next year, Pietro expects to purchase $126,000 of direct materials. Projected beginning and ending inventories for direct materials and work in process are as follows 74,000 20,000 195,000 Direct materials Work-in-Process Inventory $4,000 $3,900 Inventory $12,800 $14,800 Beginning Ending Next year, Pietro expects to produce 49,000 units and sell 48,300 units at a price of $18.00 each. Beginning inventory of finished goods is $48,500, and ending inventory of finished goods is expected to be $40,000. Total selling expense is projected at $28,500, and total administrative expense is projected at $115,500 Required: 1. Prepare an income statement in good form. Round the percent to four decimal places before converting to a percentage. For example, .88349 would be rounded to.8835 and entered as 88.351. Prepare a statement of cost of goods sold in good form. Pietro Frozen Foods, Inc. Statement of Cost of Goods Sold For the Coming Year Cost of goods manufactured Add: Beginning finished goods Cost of goods available for sale Less: Ending finished goods Cost of goods sold V Feedback Check My Work See Cornerstone 2.3 2. What if the beginning inventory of finished goods decreased by $4,500? What would be the effect on the cost of goods sold? decrease by$
Answer #1

Direct material used = 4000+126000-3900 = 126100

Total manufacturing cost = 126100+74000+20000+195000 = 415100

Cost of goods manufactured = 12800+415100-14800 = 413100

Cost of goods manufactured 413100
Beginning finished goods inventory 48500
Cost of goods available for sale 461600
Ending finished goods inventory -40000
Cost of goods sold 421600

2) cost of goods sold decrease by $4500